OpenStudy (anonymous):

19. Evaluate the sum 7 S (4 k^2+k) k=5  Possibilities: (a) 458 (b) 459 (c) 460 (d) 461 (e) 462

OpenStudy (anonymous):

the S is the Sigma sign

OpenStudy (anonymous):

why is the answer A?

OpenStudy (ash2326):

\[\sum_{k=5}^{k=7} (4k^2+k)\] As there are only three sums, we can manually evaluate \[((4 \times 5^2+5)+( 4\times 6^2+6)+(4 \times 7^2+7))\] We have \[ ( (100+5)+(144+6)+(196+7)\] \[(105)+(150)+(203)=458\]

OpenStudy (anonymous):

o so i dont have to subtract 5 from 7 and add 1 so k=1?

OpenStudy (ash2326):

no you have to sum from 5 to 7